given x = 27 ;
y = 4 ;
s = 3 ;

evaluate the given expression:

a) y*3%x

b) x + y / s
1
Expert's answer
2021-05-03T10:26:52-0400
a) Step 1. 3%x=3 mod 4 = 3
   Step 2. y=4, 4*3 =12
	Answer:  y * 3%x = 12
b) 27+(4/3)=27+1=28
where x,y,s are integer values.